# On generating mapping class groups by pseudo-Anosov elements

### v2

@@ 2. Preliminaries
and hence we have $(t_{\alpha_{2g+1}}^{2}t_{\alpha_{2g}}t_{\alpha_{2g-1}}\cdots t_{\alpha_{3}})^{g-1}=r_{\pi}\mathcal{R}r_{\pi}^{-1}t_{\alpha_{1}}\in H^{\prime}$ by $t_{\alpha_{1}}\in H^{\prime}$. We set $h=t_{\alpha_{2g+1}}^{2}t_{\alpha_{2g}}t_{\alpha_{2g-1}}\cdots t_{\alpha_{3}}$, and so $h^{g-1}$ is in $H^{\prime}$. It is easily seen that $h^{-1}(\alpha_{i})=\alpha_{i+1}$ for $i=3,4,\ldots,2g-1$, and hence $h^{-(g-1)}(\alpha_{g+1})=\alpha_{2g}$. Therefore, we have $t_{\alpha_{2g}}\in H^{\prime}$ by $t_{\alpha_{g+1}},h^{g-1}\in H^{\prime}$ and the relation $t_{h^{-(g-1)}(\alpha_{g+1})}=h^{-(g-1)}t_{\alpha_{g+1}}h^{g-1}$. Summarizing, the elements $t_{\alpha_{1}},t_{\alpha_{2}},\ldots,t_{\alpha_{2g+1}}$ are contained in $H^{\prime}$, and hence $H^{\prime}=\mathcal{H}_{g}$ for $g\geq 2$. By Theorem 2.1, $\mathcal{H}_{1}$ is generated by $t_{\alpha_{1}}$ and $t_{\alpha_{2}}$. This completes the proof. ∎
@@
In this section, Theorems 1.2 and 1.3 in the introduction. Theorem 1.2 follows from the Thorems below.
@@
For $g\geq 3$, let $G_{2}$ be the subgroup of $\mathcal{M}_{g}$ generated by $\rho_{n}$, $t_{a_{2}}\rho_{n}t_{a_{2}}^{-1}$ and $(t_{a_{3}}t_{b_{3}}t_{c_{2}})\rho_{n}(t_{a_{3}}t_{b_{3}}t_{c_{2}})^{-1}$, where $\rho_{n}$ is defined in Lemma 2.5. Then, $G_{2}=\mathcal{M}_{g}$.
Theorem 1.3 follows from the Thorems below.
@@
For $g\geq 3$, let $G_{4}$ be the subgroup of $\mathcal{M}_{g}$ generated by $\rho^{\prime}$, $t_{a_{2}}\rho^{\prime}t_{a_{2}}^{-1}$ and $(t_{a_{3}}t_{b_{3}}t_{c_{2}})\rho^{\prime}(t_{a_{3}}t_{b_{3}}t_{c_{2}})^{-1}$, where $\rho^{\prime}$ is defined in Lemma 2.6. Then, $G_{4}=\mathcal{M}_{g}$.
Lemmas 4.54.7 below are used to show Theorems 4.14.4. Recall that if a Dehn twists $t_{c}$ about a simple closed curve $c$ and an element $f$ are in a subgroup $G$ of $\mathcal{M}_{g}$, then $t_{f}(c)$ is also in $G$ by the relation $t_{f(c)}=ft_{c}f^{-1}$. Moreover, if the product $t_{c}t_{d}^{-1}$ of two Dehn twists $t_{c}$ and $t_{d}$ about two simple closed curve $c$ and $d$ and an element $f$ are in $G$, then $t_{f(c)}t_{f(d)}^{-1}$ is also in $G$ by $t_{f(c)}t_{f(d)}^{-1}=ft_{c}f^{-1}ft_{d}^{-1}f^{-1}=ft_{c}t_{d}^{-1}f^{-1}$. We will repeatedly use these facts in the proofs of Lemmas 4.54.7 and Theorems 4.14.4.
@@
for some integer $k$ such that $21\leq k\leq g-$ 2. Then, the element $t_{a_{k+2}}$ is in $G^{\prime}$.
@@
Let $d_{1}$ and $d_{2}$ be the simple closed curves in $\Sigma_{g}$ as in Figure 5. Then, it is easy to check that
$t_{b_{k+1}}t_{a_{k}}^{-1}\cdot t_{c_{k}}t_{a_{k}}^{-1}\cdot t_{a_{k}}t_{a_{k+1}}^{-1}\cdot t_{c_{k+1}}t_{a_{k}}^{-1}(b_{k+1},a_{k})$ $=(d_{1},a_{k})$
$t_{b_{k+2}}t_{a_{k}}^{-1}\cdot t_{c_{k+1}}t_{a_{k}}^{-1}\cdot t_{a_{k+2}}t_{a_{k}}^{-1}\cdot t_{b_{k+2}}t_{a_{k}}^{-1}(d_{1},a_{k})$ $=(d_{2},a_{k}).$
When we set $\phi_{1}=t_{b_{k+1}}t_{a_{k}}^{-1}\cdot t_{c_{k}}t_{a_{k}}^{-1}\cdot t_{a_{k}}t_{a_{k+1}}^{-1}\cdot t_{c_{k+1}}t_{a_{k}}^{-1}$ and $\phi_{2}=t_{b_{k+2}}t_{a_{k}}^{-1}\cdot t_{c_{k+1}}t_{a_{k}}^{-1}\cdot t_{a_{k+2}}t_{a_{k}}^{-1}\cdot t_{b_{k+2}}t_{a_{k}}^{-1}$, by $\phi_{1},\phi_{2},t_{b_{k+1}}t_{a_{k}}^{-1}\in G^{\prime}$ we see that
$t_{d_{1}}t_{a_{k}}^{-1}\in G^{\prime},$
$t_{d_{2}}t_{a_{k}}^{-1}=\phi_{2}\cdot t_{d_{1}}t_{a_{k}}^{-1}\cdot\phi_{2}^{-1}\in G^{\prime}.$
Moreover, by $t_{a_{k}}t_{c_{k}}^{-1}\in G^{\prime}$, we have $t_{d_{2}}t_{c_{k}}^{-1}=t_{d_{2}}t_{a_{k}}^{-1}\cdot t_{a_{k}}t_{c_{k}}^{-1}\in G^{\prime}$. Therefore, by the lantern relation $t_{a_{k}}t_{c_{k}}t_{c_{k+1}}t_{a_{k+2}}=t_{a_{k+1}}t_{d_{1}}t_{d_{2}}$, we have
$t_{a_{k+2}}=t_{a_{k+1}}t_{c_{k+1}}^{-1}\cdot t_{d_{1}}t_{a_{k}}^{-1}\cdot t_{d_{2}}t_{c_{k}}^{-1}\in G^{\prime}.$
@@
By $f,t_{a_{k}}t_{a_{k+1}}^{-1},t_{b_{k}}t_{b_{k+1}}^{-1}\in G^{\prime}$, the following holds:
$t_{a_{k+1}}t_{a_{k+2}}^{-1}\in G^{\prime},$
$t_{b_{k+1}}t_{b_{k+2}}^{-1}\in G^{\prime}.$
Since it is easy to check that $t_{a_{k}}t_{a_{k+1}}^{-1}\cdot t_{b_{k}}t_{b_{k+1}}^{-1}(a_{k+1},a_{k+2})=(b_{k+1},a_{k+2})$, by $t_{a_{k}}t_{a_{k+1}}^{-1},t_{b_{k}}t_{b_{k+1}}^{-1},t_{a_{k+1}}t_{a_{k+2}}^{-1}\in G^{\prime}$, we have
$t_{b_{k+1}}t_{a_{k+2}}^{-1}\in G^{\prime},$
$t_{b_{k}}t_{a_{k+2}}^{-1}=t_{b_{k}}t_{b_{k+1}}^{-1}\cdot t_{b_{k+1}}t_{a_{k+2}}^{-1}\in G^{\prime}.$
Similarly, since it is easily seen that $t_{b_{k}}t_{a_{k+2}}^{-1}\cdot t_{c_{k}}t_{c_{k+1}}^{-1}(b_{k},a_{k+2})=(c_{k},a_{k+2})$, by $t_{b_{k}}t_{a_{k+2}}^{-1},t_{c_{k}}t_{c_{k+1}}^{-1}\in G^{\prime}$, we have
$t_{c_{k}}t_{a_{k+2}}^{-1}\in G^{\prime}.$
We pick up elements in the above arguments to prove Lemma 4.6:
$t_{a_{k}}t_{a_{k+1}}^{-1},\ t_{a_{k+1}}t_{a_{k+2}}^{-1},\ t_{b_{k+1}}t_{a_{k+2}}^{-1},\ t_{c_{k}}t_{a_{k+2}}^{-1}\in G^{\prime}.$
Using these elements, we have
$t_{a_{k}}t_{a_{k+2}}^{-1}$ $=t_{a_{k}}t_{a_{k+1}}^{-1}\cdot t_{a_{k+1}}t_{a_{k+2}}^{-1}\in G^{\prime},$
$t_{a_{k}}t_{b_{k+1}}^{-1}$ $=t_{a_{k}}t_{a_{k+1}}^{-1}\cdot t_{a_{k+1}}t_{a_{k+2}}^{-1}\cdot(t_{b_{k+1}}t_{a_{k+2}}^{-1})^{-1}\in G^{\prime},$
$t_{a_{k}}t_{c_{k}}^{-1}$ $=t_{a_{k}}t_{a_{k+1}}^{-1}\cdot t_{a_{k+1}}t_{a_{k+2}}^{-1}\cdot(t_{c_{k}}t_{a_{k+2}}^{-1})^{-1}\in G^{\prime},$
$t_{a_{k+1}}t_{c_{k+1}}^{-1}$ $=t_{f(a_{k})}t_{f(c_{k})}^{-1}=ft_{a_{k}}t_{c_{k}}^{-1}f^{-1}\in G^{\prime},$
$t_{a_{k}}t_{c_{k+1}}^{-1}$ $=t_{a_{k}}t_{a_{k+1}}^{-1}\cdot t_{a_{k+1}}t_{c_{k+1}}^{-1}\in G^{\prime}.$
Moreover, since we have $t_{a_{k+1}}t_{b_{k+2}}^{-1}\in G^{\prime}$ by the assumption that $f,t_{a_{k}}t_{b_{k+1}}^{-1}\in G^{\prime}$, the element $t_{a_{k}}t_{b_{k+2}}^{-1}$ is in $G^{\prime}$ by $t_{a_{k}}t_{b_{k+2}}^{-1}=t_{a_{k}}t_{a_{k+1}}^{-1}\cdot t_{a_{k+1}}t_{b_{k+2}}^{-1}$. Summarizing, we have
$t_{a_{k}}t_{a_{k+1}}^{-1},\ t_{a_{k}}t_{a_{k+2}}^{-1},\ t_{a_{k}}t_{b_{k+1}}^{-1},\ t_{a_{k}}t_{c_{k}}^{-1},\ t_{a_{k+1}}t_{c_{k+1}}^{-1},\ t_{a_{k}}t_{c_{k+1}}^{-1},\ t_{a_{k}}t_{b_{k+2}}^{-1}\in G^{\prime},$
and we see that $t_{a_{k+2}}\in G^{\prime}$ from Lemma 4.5. By $t_{a_{k+2}}\in G^{\prime}$ and the assumption that $f^{2}(a_{k})=a_{k+2}$, the element $t_{a_{k}}$ is contained in $G^{\prime}$. By $t_{a_{k}}t_{b_{k+1}}^{-1},t_{a_{k}}t_{c_{k+1}}^{-1}\in G^{\prime}$, we have $t_{b_{k+1}},t_{c_{k+1}}\in G^{\prime}$. Similarly, by $t_{b_{k+1}},t_{c_{k+1}}\in G^{\prime}$ and the assumption that $f^{-1}(b_{k+1},c_{k+1})=(b_{k},c_{k})$, we have $t_{b_{k}},t_{c_{k}}\in G^{\prime}$, and the lemma follows. ∎
@@
Note that $\Phi_{k}\in G^{\prime\prime}$. By the assumption of $f$, we see that $\Phi_{k+1}=t_{f(b_{k})}t_{f(a_{k+2})}^{-1}t_{f(c_{k+3})}\cdot t_{f(c_{k+4})}^{-1}t_{f(a_{k+3})}t_{f(b_{k+1})}^{-1}=f\Phi_{k}f^{-1}\in G^{\prime\prime}$. Since it is easy to check that
@@
we have
$t_{b_{k+1}}t_{a_{k+3}}^{-1}t_{c_{k+4}}\cdot t_{c_{k+5}}^{-1}t_{a_{k+4}}t_{a_{k+2}}^{-1}=(\Phi_{k+1}\Phi_{k})\Phi_{k+1}(\Phi_{k+1}\Phi_{k})^{-1}\in G^{\prime\prime}.$
Then, we obtain the following:
@@
By $t_{a_{k+2}}t_{b_{k+2}}^{-1}\in G^{\prime\prime}$ and the assumption of $f$, we have
(2) $t_{a_{k+3}}t_{b_{k+3}}^{-1}\in G^{\prime\prime},$
@@
Here, since the curves $b_{k+1},a_{k+3},c_{k+4},c_{k+5},a_{k+4}$ are disjoint from $a_{k+2}$ and $b_{k+2}$, we have $t_{b_{k+2}}^{-1}t_{a_{k+2}}=\Phi_{k+1}(t_{b_{k+1}}t_{a_{k+3}}^{-1}t_{c_{k+4}}\cdot t_{c_{k+5}}^{-1}t_{a_{k+4}}t_{a_{k+2}}^{-1})^{-1}\in G^{\prime\prime}$. By similar arguments to (2) and (3), we obtain the following:
(4) $t_{b_{k+3}}^{-1}t_{a_{k+3}}\in G^{\prime\prime},$
@@
By (4), we see that $\Phi_{k}t_{a_{k+3}}^{-1}t_{b_{k+3}}=\Phi_{k}(t_{b_{k+3}}^{-1}t_{a_{k+3}})^{-1}\in G^{\prime\prime}$. Since $b_{k+1}$ is disjoint from $a_{k+3}$ and $b_{k+3}$, we have
@@
$=t_{b_{k}}t_{a_{k+2}}^{-1}t_{c_{k+3}}\cdot t_{c_{k+4}}^{-1}t_{b_{k+3}}t_{b_{k+1}}^{-1}\ (\in G^{\prime\prime}).$
From this, we see that $(\Phi_{k}t_{a_{k+3}}^{-1}t_{b_{k+3}})^{-1}(a_{k+3},b_{k+3})=(t_{b_{k+3}}^{-1}(a_{k+3}),c_{k+3})$, and therefore, by $t_{a_{k+3}}t_{b_{k+3}}^{-1},(\Phi_{k}t_{a_{k+3}}^{-1}t_{b_{k+3}})^{-1}\in G^{\prime\prime}$, the following holds:
$t_{b_{k+3}}^{-1}t_{a_{k+3}}t_{b_{k+3}}t_{c_{k+3}}^{-1}=t_{t_{b_{k+3}}^{-1}(a_{k+3})}t_{c_{k+3}}^{-1}\in G^{\prime\prime}.$
By (4), we have
@@
By (5) and $\Phi_{k},\Phi_{k+1}\in G^{\prime\prime}$, we see that $\Phi_{k}\Phi_{k+1}t_{a_{k+4}}t_{b_{k+4}}^{-1}=\Phi_{k}\Phi_{k+1}(t_{b_{k+4}}^{-1}t_{a_{k+4}})^{-1}\in G^{\prime\prime}$. Since $b_{k+2}$ are disjoint from $a_{k+4}$ and $b_{k+4}$, we have
$\Phi_{k}\Phi_{k+1}t_{a_{k+4}}t_{b_{k+4}}^{-1}$ $=t_{b_{k}}t_{a_{k+2}}^{-1}t_{c_{k+3}}\cdot t_{c_{k+5}}^{-1}t_{a_{k+4}}t_{b_{k+2}}^{-1}\cdot(t_{b_{k+4}}^{-1}t_{a_{k+4}})^{-1}$
$=t_{b_{k}}t_{a_{k+2}}^{-1}t_{c_{k+3}}\cdot t_{c_{k+5}}^{-1}t_{b_{k+4}}t_{b_{k+2}}^{-1}\ (\in G^{\prime\prime}).$
From this, we see that $(\Phi_{k}\Phi_{k+1}t_{a_{k+4}}t_{b_{k+4}}^{-1})^{-1}(a_{k+4},b_{k+4})=(t_{b_{k+4}}^{-1}(a_{k+4}),c_{k+3})$, and therefore, by $t_{a_{k+4}}t_{b_{k+4}}^{-1},(\Phi_{k}\Phi_{k+1}t_{a_{k+4}}t_{b_{k+4}}^{-1})^{-1}\in G^{\prime\prime}$, the following holds:
$t_{b_{k+4}}^{-1}t_{a_{k+4}}t_{b_{k+4}}t_{c_{k+3}}^{-1}=t_{t_{b_{k+4}}^{-1}(a_{k+4})}t_{c_{k+3}}^{-1}\in G^{\prime\prime}.$
By $t_{b_{k+4}}^{-1}t_{a_{k+4}}t_{b_{k+4}}t_{c_{k+3}}^{-1}\in G^{\prime\prime}$, (5) and the assumption that $f^{-1}(b_{k+4},c_{k+3})=(b_{k+3},c_{k+2})$, we have
(7) $t_{b_{k+4}}t_{c_{k+3}}^{-1}\in G^{\prime\prime},$

### v3

@@ 2. Preliminaries
and hence $(t_{\alpha_{2g+1}}^{2}t_{\alpha_{2g}}t_{\alpha_{2g-1}}\cdots t_{\alpha_{3}})^{g-1}=r_{\pi}\mathcal{R}r_{\pi}^{-1}t_{\alpha_{1}}$ is in $H^{\prime}$ since $t_{\alpha_{1}}$ is in $H^{\prime}$. When we set $h=t_{\alpha_{2g+1}}^{2}t_{\alpha_{2g}}t_{\alpha_{2g-1}}\cdots t_{\alpha_{3}}$, $h^{g-1}$ is in $H^{\prime}$. It is easily seen that $h^{-1}(\alpha_{i})=\alpha_{i+1}$ for $i=3,4,\ldots,2g-1$, and hence $h^{-(g-1)}(\alpha_{g+1})=\alpha_{2g}$. Therefore, $t_{\alpha_{2g}}$ is in $H^{\prime}$ since $t_{\alpha_{g+1}}$ and $h^{g-1}$ are contained in $H^{\prime}$. Summarizing, the elements $t_{\alpha_{1}},t_{\alpha_{2}},\ldots,t_{\alpha_{2g+1}}$ are contained in $H^{\prime}$, and hence $H^{\prime}=\mathcal{H}_{g}$ for $g\geq 2$. By Theorem 2.1, $\mathcal{H}_{1}$ is generated by $t_{\alpha_{1}}$ and $t_{\alpha_{2}}$. This completes the proof. ∎
@@
In this section, we prove Theorems 1.2 and 1.3 in the introduction. Theorem 1.2 (resp. 1.3) follows from Theorems 4.1 and 4.2 (resp. 4.3 and 4.4) below.
@@
Let $g\geq 3$, and let $G_{2}$ be the subgroup of $\mathcal{M}_{g}$ generated by $\rho_{n}$, $t_{a_{2}}\rho_{n}t_{a_{2}}^{-1}$ and $(t_{a_{3}}t_{b_{3}}t_{c_{2}})\rho_{n}(t_{a_{3}}t_{b_{3}}t_{c_{2}})^{-1}$, where $\rho_{n}$ is defined in Lemma 2.5. Then, $G_{2}=\mathcal{M}_{g}$.
@@
Let $g\geq 3$, and let $G_{4}$ be the subgroup of $\mathcal{M}_{g}$ generated by $\rho^{\prime}$, $t_{a_{2}}\rho^{\prime}t_{a_{2}}^{-1}$ and $(t_{a_{3}}t_{b_{3}}t_{c_{2}})\rho^{\prime}(t_{a_{3}}t_{b_{3}}t_{c_{2}})^{-1}$, where $\rho^{\prime}$ is defined in Lemma 2.6. Then, $G_{4}=\mathcal{M}_{g}$.
Lemmas 4.54.7 below are used to show Theorems 4.14.4. Recall that if a Dehn twist $t_{c}$ and an element $f$ are in a subgroup $G$ of $\mathcal{M}_{g}$, then $t_{f}(c)$ is also in $G$ by $t_{f(c)}=ft_{c}f^{-1}$. Therefore, if an element $h$ and a product $t_{c}t_{d}^{-1}$ of two Dehn twists $t_{c}$ and $t_{d}$ are in $G$, then $t_{h(c)}t_{h(d)}^{-1}$ is also in $G$. Let us consider the simple closed curves in Figure 2. Note that $a_{i}$ (resp. $b_{i}$) and $b_{i}$ (resp. $c_{i}$) intersects transversely at exactly one point for $i=1,2,\ldots,g$, and that $b_{i}$ (resp. $b_{1}$) and $c_{i-1}$ (resp. $c_{g}$) intersects transversely at exactly one point for $i=2,3,\ldots,g$. All other pairs of curves are disjoint. We will repeatedly use these facts in the proofs of Lemmas 4.54.7 and Theorems 4.14.4.
@@
for some integer $k$ such that $1\leq k\leq g-2$. Then, the element $t_{a_{k+2}}$ is in $G^{\prime}$.
@@
Let $d_{1}$ and $d_{2}$ be the simple closed curves in $\Sigma_{g}$ as in Figure 5. We set $\phi_{1}=t_{b_{k+1}}t_{a_{k}}^{-1}\cdot t_{c_{k}}t_{a_{k}}^{-1}\cdot t_{a_{k}}t_{a_{k+1}}^{-1}\cdot t_{c_{k+1}}t_{a_{k}}^{-1}$ and $\phi_{2}=t_{b_{k+2}}t_{a_{k}}^{-1}\cdot t_{c_{k+1}}t_{a_{k}}^{-1}\cdot t_{a_{k+2}}t_{a_{k}}^{-1}\cdot t_{b_{k+2}}t_{a_{k}}^{-1}$. Then, it is easy to check that $\phi_{1}(b_{k+1},a_{k})=(d_{1},a_{k})$ and $\phi_{2}(d_{1},a_{k})=(d_{2},a_{k})$. Therefore, $t_{d_{1}}t_{a_{k}}^{-1}=(\phi_{1}t_{a_{k}}t_{b_{k+1}}^{-1}\phi_{1}^{-1})^{-1}$ and $t_{d_{2}}t_{a_{k}}^{-1}=\phi_{2}\cdot t_{d_{1}}t_{a_{k}}^{-1}\cdot\phi_{2}^{-1}$ are in $G^{\prime}$ since $\phi_{1},\phi_{2}$ and $t_{a_{k}}t_{b_{k+1}}^{-1}$ are in $G^{\prime}$. Moreover, $t_{d_{2}}t_{c_{k}}^{-1}=t_{d_{2}}t_{a_{k}}^{-1}\cdot t_{a_{k}}t_{c_{k}}^{-1}$ is in $G^{\prime}$ since $t_{a_{k}}t_{c_{k}}^{-1}$ is in $G^{\prime}$. Therefore, by the lantern relation $t_{a_{k}}t_{c_{k}}t_{c_{k+1}}t_{a_{k+2}}=t_{a_{k+1}}t_{d_{1}}t_{d_{2}}$, we obtain that $t_{a_{k+2}}=t_{a_{k+1}}t_{c_{k+1}}^{-1}\cdot t_{d_{1}}t_{a_{k}}^{-1}\cdot t_{d_{2}}t_{c_{k}}^{-1}$ is in $G^{\prime}$. ∎
@@
Since $f,t_{a_{k}}t_{a_{k+1}}^{-1}$ and $t_{b_{k}}t_{b_{k+1}}^{-1}$ are in $G^{\prime}$, we see that $t_{a_{k+1}}t_{a_{k+2}}^{-1}$ and $t_{b_{k+1}}t_{b_{k+2}}^{-1}$ are in $G^{\prime}$. Note that it is easy to check that $t_{a_{k}}t_{a_{k+1}}^{-1}\cdot t_{b_{k}}t_{b_{k+1}}^{-1}(a_{k+1},a_{k+2})=(b_{k+1},a_{k+2})$. Since $t_{a_{k}}t_{a_{k+1}}^{-1},t_{b_{k}}t_{b_{k+1}}^{-1}$ and $t_{a_{k+1}}t_{a_{k+2}}^{-1}$ are contained in $G^{\prime}$, we see that $t_{b_{k+1}}t_{a_{k+2}}^{-1}$ and $t_{b_{k}}t_{a_{k+2}}^{-1}=t_{b_{k}}t_{b_{k+1}}^{-1}\cdot t_{b_{k+1}}t_{a_{k+2}}^{-1}$ are also contained in $G^{\prime}$. Similarly, since $t_{b_{k}}t_{a_{k+2}}^{-1}$ and $t_{c_{k}}t_{c_{k+1}}^{-1}$ are in $G^{\prime}$ and it is easily seen that $t_{b_{k}}t_{a_{k+2}}^{-1}\cdot t_{c_{k}}t_{c_{k+1}}^{-1}(b_{k},a_{k+2})=(c_{k},a_{k+2})$, we obtain that $t_{c_{k}}t_{a_{k+2}}^{-1}$ is in $G^{\prime}$. By the above argument, we showed that the following elements are in $G^{\prime}$:
$t_{a_{k}}t_{a_{k+1}}^{-1},\ t_{a_{k+1}}t_{a_{k+2}}^{-1},\ t_{b_{k+1}}t_{a_{k+2}}^{-1},\ t_{c_{k}}t_{a_{k+2}}^{-1}.$
Using these elements, we see that the elements
$t_{a_{k}}t_{a_{k+2}}^{-1}$ $=t_{a_{k}}t_{a_{k+1}}^{-1}\cdot t_{a_{k+1}}t_{a_{k+2}}^{-1},$
$t_{a_{k}}t_{b_{k+1}}^{-1}$ $=t_{a_{k}}t_{a_{k+1}}^{-1}\cdot t_{a_{k+1}}t_{a_{k+2}}^{-1}\cdot(t_{b_{k+1}}t_{a_{k+2}}^{-1})^{-1},$
$t_{a_{k}}t_{c_{k}}^{-1}$ $=t_{a_{k}}t_{a_{k+1}}^{-1}\cdot t_{a_{k+1}}t_{a_{k+2}}^{-1}\cdot(t_{c_{k}}t_{a_{k+2}}^{-1})^{-1},$
$t_{a_{k+1}}t_{c_{k+1}}^{-1}$ $=t_{f(a_{k})}t_{f(c_{k})}^{-1}=ft_{a_{k}}t_{c_{k}}^{-1}f^{-1},$
$t_{a_{k}}t_{c_{k+1}}^{-1}$ $=t_{a_{k}}t_{a_{k+1}}^{-1}\cdot t_{a_{k+1}}t_{c_{k+1}}^{-1},$
$t_{a_{k+1}}t_{b_{k+2}}^{-1}$ $=t_{f(a_{k})}t_{f(b_{k+1})}^{-1}=ft_{a_{k}}t_{b_{k+1}}f^{-1},\ \ \mathrm{and}$
$t_{a_{k}}t_{b_{k+2}}^{-1}$ $=t_{a_{k}}t_{a_{k+1}}^{-1}\cdot t_{a_{k+1}}t_{b_{k+2}}^{-1}$
are contained in $G^{\prime}$. As a consequence, we verified that $G^{\prime}$ satisfies the condition on Lemma  4.5, that is
$t_{a_{k}}t_{a_{k+1}}^{-1},\ t_{a_{k}}t_{a_{k+2}}^{-1},\ t_{a_{k}}t_{b_{k+1}}^{-1},\ t_{a_{k}}t_{c_{k}}^{-1},\ t_{a_{k+1}}t_{c_{k+1}}^{-1},\ t_{a_{k}}t_{c_{k+1}}^{-1},\ t_{a_{k}}t_{b_{k+2}}^{-1},$
are elements of $G^{\prime}$. Then we see that $t_{a_{k+2}}$ is in $G^{\prime}$ by applying Lemma 4.5. Moreover, the element $t_{a_{k}}$ is contained in $G^{\prime}$ by the assumption that $f^{2}(a_{k})=a_{k+2}$. Therefore, $t_{b_{k+1}}$ and $t_{c_{k+1}}$ are contained in $G^{\prime}$ since $G^{\prime}$ contains $t_{a_{k}}t_{b_{k+1}}^{-1}$ and $t_{a_{k}}t_{c_{k+1}}^{-1}$. Similarly, by the assumption that $f^{-1}(b_{k+1},c_{k+1})=(b_{k},c_{k})$, $t_{b_{k}}$ and $t_{c_{k}}$ are contained in $G^{\prime}$ since $t_{b_{k+1}}$, $t_{c_{k+1}}$ are in $G^{\prime}$, and the lemma follows. ∎
@@
Note that $\Phi_{k}$ is in $G^{\prime\prime}$. Then, $\Phi_{k+1}=f\Phi_{k}f^{-1}$ is also $G^{\prime\prime}$ by the assumption on $f$. Since it is easy to check that
@@
we see that the element
$t_{b_{k+1}}t_{a_{k+3}}^{-1}t_{c_{k+4}}\cdot t_{c_{k+5}}^{-1}t_{a_{k+4}}t_{a_{k+2}}^{-1}=(\Phi_{k+1}\Phi_{k})\Phi_{k+1}(\Phi_{k+1}\Phi_{k})^{-1}$
is contained in $G^{\prime\prime}$. Then, we obtain the following:
@@
Moreover, from the assumption on $f$, the following holds:
(2) $t_{a_{k+3}}t_{b_{k+3}}^{-1}\in G^{\prime\prime}\ and$
@@
Here, we find that $t_{b_{k+2}}^{-1}t_{a_{k+2}}=\Phi_{k+1}(t_{b_{k+1}}t_{a_{k+3}}^{-1}t_{c_{k+4}}\cdot t_{c_{k+5}}^{-1}t_{a_{k+4}}t_{a_{k+2}}^{-1})^{-1}$, and hence $t_{b_{k+2}}^{-1}t_{a_{k+2}}$ is in $G^{\prime\prime}$. By similar arguments to (2) and (3), we obtain the following:
(4) $t_{b_{k+3}}^{-1}t_{a_{k+3}}\in G^{\prime\prime},\ \ \mathrm{and}$
@@
By (4), we see that $\Phi_{k}t_{a_{k+3}}^{-1}t_{b_{k+3}}=\Phi_{k}(t_{b_{k+3}}^{-1}t_{a_{k+3}})^{-1}$ is in $G^{\prime\prime}$. Moreover, we have
@@
$=t_{b_{k}}t_{a_{k+2}}^{-1}t_{c_{k+3}}\cdot t_{c_{k+4}}^{-1}t_{b_{k+3}}t_{b_{k+1}}^{-1}.$
This gives $(\Phi_{k}t_{a_{k+3}}^{-1}t_{b_{k+3}})^{-1}(a_{k+3},b_{k+3})=(t_{b_{k+3}}^{-1}(a_{k+3}),c_{k+3})$. Therefore, since $t_{a_{k+3}}t_{b_{k+3}}^{-1}$ and $(\Phi_{k}t_{a_{k+3}}^{-1}t_{b_{k+3}})^{-1}$ are in $G^{\prime\prime}$, the element $t_{t_{b_{k+3}}^{-1}(a_{k+3})}t_{c_{k+3}}^{-1}$ is in $G^{\prime\prime}$. That is, $t_{b_{k+3}}^{-1}t_{a_{k+3}}t_{b_{k+3}}t_{c_{k+3}}^{-1}$ is in $G^{\prime\prime}$. By (4), we have the following:
@@
Since $\Phi_{k}$ and $\Phi_{k+1}$ are contained in $G^{\prime\prime}$, by (5) we see that $\Phi_{k}\Phi_{k+1}t_{a_{k+4}}^{-1}t_{b_{k+4}}=\Phi_{k}\Phi_{k+1}(t_{b_{k+4}}^{-1}t_{a_{k+4}})^{-1}$ is in $G^{\prime\prime}$. Therefore, we obtain that
$\Phi_{k}\Phi_{k+1}t_{a_{k+4}}^{-1}t_{b_{k+4}}$ $=t_{b_{k}}t_{a_{k+2}}^{-1}t_{c_{k+3}}\cdot t_{c_{k+5}}^{-1}t_{a_{k+4}}t_{b_{k+2}}^{-1}\cdot t_{a_{k+4}}^{-1}t_{b_{k+4}}$
$=t_{b_{k}}t_{a_{k+2}}^{-1}t_{c_{k+3}}\cdot t_{c_{k+5}}^{-1}t_{b_{k+4}}t_{b_{k+2}}^{-1}$
is in $G^{\prime\prime}$. This gives $(\Phi_{k}\Phi_{k+1}t_{a_{k+4}}^{-1}t_{b_{k+4}})^{-1}(a_{k+4},b_{k+4})=(t_{b_{k+4}}^{-1}(a_{k+4}),c_{k+3})$, and hence we see that $t_{t_{b_{k+4}}^{-1}(a_{k+4})}t_{c_{k+3}}^{-1}$ is in $G^{\prime\prime}$ by (3). That is, $t_{b_{k+4}}^{-1}t_{a_{k+4}}t_{b_{k+4}}t_{c_{k+3}}^{-1}$ is in $G^{\prime\prime}$. Therefore, by (5) and the assumption that $f^{-1}(b_{k+4},c_{k+3})=(b_{k+3},c_{k+2})$, the following holds:
(7) $t_{b_{k+4}}t_{c_{k+3}}^{-1}\in G^{\prime\prime},\ \ \mathrm{and}$